g quegtioHs la the text box below or by uploading a word document. Show work on all calcu lations. This is new for you all so I want to be able to help you through any mistakes your that happen ) 1. A patient is receiving 1328 mL of a 2 in 1 solution of 4AA/D25W. Calculate the kcals provided by amino 2. Using the same solution from question 1, calculate the estimated osmolarity of this volume of solution. 3. A patient weighing 82 kg is receiving a 20% lipid solution at 21 mL per hour. Is this rate of lipid infusion 4. A patient is determined to need a NPC:N of 100:1. The patient is receiving 1800 kcal total from their acids and the kcals provided by dextrose in this volume of solution. Based on the osmolarity you calculated, which access would be safest for administration? appropriate when considering the max rate of lipid infusion that is considered safe? solutions. The patient is receiving 1500 mL of a 4%AA solution. Is this volume meeting the the estimated NPC:N needs?
Answer #1

1. Total volume of solution = 1328 ml

Solution containing 4AA / D25W

In grams 4AA= 40gm

In grams dextrose = 250gm

1gm of each protein and carbohydrates provide 4Kcal

So, 40gm of protein provides = 40×4=160 Kcal

250gm of dextrose provides = 250 × 4 = 1000Kcal

2 . To calculate the osmolarity of the solution :-

Multiply grams of dextrose per litre by 5

Multiply grams of protein per litre by 10

Add both of them

Then add 300- 400 mOsm/ L to it as vitamins and minerals contribute this much in a solution .

Now applying them ,

D25W = 250g dextrose /l

4AA= 40g AA/l

250 × 5= 1250 mOsm

40×10 = 400 mOsm

Add them = 1650 mOsm

Then add vitamins and minerals (300-400) to it

= 1650+300 to 400

= 1950 to 2050mOsm/L